# Constraining the Stellar-to-Halo Mass Relation with Galaxy Clustering and Weak Lensing from DES Year 3 Data

### v2

@@ Appendix E Diagnostic tests
  • Random points: The mean tangential shear around random points tests the importance of geometrical and mask effects in the signal. Although our estimator of galaxy-galaxy lensing includes the subtraction of tangential shear measurement around random points, it is useful to check that this correction is small, especially for the bins with the highest signal. For this work we use a number of random points that is $30$ times the number of lenses in each bin. We note here, however, that we have tested different numbers of random points, both higher and lower, and we have verified that this number is sufficiently large for our purposes. See also Prat et al. (2022) for a discussion on random-point subtraction and its effects on the measured shear. The effect from including the random-point subtraction is shown on Figure 18, where we can see that it is a small effect, and within the error bars in all bins, as also indicated by the reduced $\chi^{2}$ values we obtain.
@@
  • Boost factors: Galaxies are clustered, and the number of sources around a lens can be larger than what we would expect from a uniform distribution if the lens and the source redshift distribution overlap. This is usually quantified by the boost factor (Sheldon et al., 2004), $B(\theta)$, as estimator quantifying the excess in the number of sources around a lens with respect to randoms. This is a test only relevant for $\gamma_{t}$ and the results are plotted in Figure 18. As can be seen from the plot, the contribution from this effect can be large at small scales, especially when the bins are more overlapped in redshift. In our analysis we take the boost factors into account by correcting for it before carrying out the model fit (Equation 5.1). We also use the boost factors to guide our decision in the selection of bins and scales to use in the final analysis (Section 5.4).
@@ Appendix F Constraint on parameters
(1, 1) $0.427^{\small+0.144}_{\small-0.165}$ $1.378^{\small+0.079}_{\small-0.082}$ $14.077^{\small+2.551}_{\small-3.090}$ $0.202^{\small+0.073}_{\small-0.061}$ $9.411^{\small+3.051}_{\small-2.936}$ $0.419^{\small+0.228}_{\small-0.244}$ $11.624^{\small+0.101}_{\small-0.104}$ $-1.068^{\small+0.172}_{\small-0.189}$ $-2.007^{\small+0.329}_{\small-0.158}$ $0.208^{\small+0.192}_{\small-0.123}$ $5.883^{\small+0.473}_{\small-0.743}$
(1, 2) $0.329^{\small+0.069}_{\small-0.165}$ $0.628^{\small+0.046}_{\small-0.088}$ $7.483^{\small+6.117}_{\small-2.162}$ $1.739^{\small+0.205}_{\small-0.253}$ $10.133^{\small+3.321}_{\small-2.296}$ $0.141^{\small+0.380}_{\small-0.351}$ $11.168^{\small+0.157}_{\small-0.186}$ $-1.736^{\small+0.095}_{\small-0.101}$ $-1.940^{\small+0.329}_{\small-0.226}$ $0.555^{\small+0.198}_{\small-0.116}$ $3.807^{\small+0.576}_{\small-0.589}$
(1, 3) $0.142^{\small+0.127}_{\small-0.101}$ $0.791^{\small+0.046}_{\small-0.059}$ $7.552^{\small+1.024}_{\small-1.133}$ $1.999^{\small+0.280}_{\small-0.271}$ $7.295^{\small+3.541}_{\small-4.336}$ $0.767^{\small+0.278}_{\small-0.710}$ $11.287^{\small+0.094}_{\small-0.212}$ $-1.700^{\small+0.093}_{\small-0.089}$ $-2.016^{\small+0.262}_{\small-0.251}$ $0.679^{\small+0.131}_{\small-0.086}$ $3.874^{\small+0.658}_{\small-0.222}$
(1, 4) $0.146^{\small+0.015}_{\small-0.055}$ $1.087^{\small+0.004}_{\small-0.050}$ $8.193^{\small+0.027}_{\small-2.736}$ $1.554^{\small+0.290}_{\small-0.320}$ $5.592^{\small+1.800}_{\small-1.173}$ $-0.284^{\small+0.485}_{\small-0.269}$ $11.147^{\small+0.437}_{\small-0.071}$ $-1.893^{\small+0.082}_{\small-0.035}$ $-1.977^{\small+0.589}_{\small-0.022}$ $0.855^{\small+0.089}_{\small-0.103}$ $5.414^{\small+0.130}_{\small-0.615}$
(2, 1) $0.492^{\small+0.060}_{\small-0.099}$ $1.232^{\small+0.027}_{\small-0.030}$ $18.918^{\small+1.351}_{\small-1.817}$ $1.289^{\small+0.129}_{\small-0.129}$ $0.963^{\small+3.079}_{\small-2.076}$ $1.446^{\small+0.443}_{\small-0.654}$ $10.519^{\small+0.114}_{\small-0.095}$ $-1.968^{\small+0.088}_{\small-0.065}$ $-1.220^{\small+0.308}_{\small-0.277}$ $0.796^{\small+0.062}_{\small-0.093}$ $4.652^{\small+0.260}_{\small-0.351}$
(2, 2) $0.414^{\small+0.092}_{\small-0.129}$ $1.230^{\small+0.053}_{\small-0.052}$ $9.170^{\small+2.976}_{\small-3.270}$ $1.147^{\small+0.200}_{\small-0.245}$ $0.864^{\small+3.844}_{\small-3.268}$ $1.524^{\small+0.620}_{\small-0.677}$ $10.609^{\small+0.156}_{\small-0.141}$ $-1.744^{\small+0.125}_{\small-0.124}$ $-1.571^{\small+0.288}_{\small-0.302}$ $0.623^{\small+0.130}_{\small-0.126}$ $3.948^{\small+0.522}_{\small-0.554}$
(2, 3) $0.442^{\small+0.109}_{\small-0.125}$ $1.232^{\small+0.051}_{\small-0.055}$ $8.058^{\small+2.303}_{\small-2.142}$ $1.541^{\small+0.182}_{\small-0.245}$ $1.034^{\small+2.307}_{\small-1.468}$ $0.348^{\small+0.678}_{\small-0.666}$ $11.113^{\small+0.180}_{\small-0.180}$ $-1.864^{\small+0.157}_{\small-0.120}$ $-1.577^{\small+0.304}_{\small-0.312}$ $0.635^{\small+0.102}_{\small-0.126}$ $4.322^{\small+0.527}_{\small-0.587}$
(2, 4) $0.487^{\small+0.250}_{\small-0.138}$ $1.173^{\small+0.037}_{\small-0.055}$ $4.053^{\small+1.528}_{\small-1.217}$ $1.560^{\small+0.360}_{\small-0.138}$ $0.300^{\small+1.785}_{\small-0.993}$ $0.710^{\small+0.371}_{\small-0.524}$ $11.857^{\small+0.385}_{\small-0.473}$ $-1.870^{\small+0.228}_{\small-0.095}$ $-2.056^{\small+0.426}_{\small-0.295}$ $0.599^{\small+0.170}_{\small-0.052}$ $3.740^{\small+0.932}_{\small-0.393}$
(3, 1) $0.236^{\small+0.111}_{\small-0.105}$ $1.165^{\small+0.063}_{\small-0.066}$ $7.050^{\small+2.638}_{\small-2.288}$ $1.004^{\small+0.197}_{\small-0.213}$ $2.622^{\small+3.662}_{\small-2.779}$ $1.527^{\small+0.683}_{\small-0.717}$ $10.746^{\small+0.164}_{\small-0.154}$ $-1.545^{\small+0.140}_{\small-0.151}$ $-1.730^{\small+0.312}_{\small-0.279}$ $0.484^{\small+0.152}_{\small-0.121}$ $3.441^{\small+0.578}_{\small-0.466}$
(3, 2) $0.259^{\small+0.128}_{\small-0.091}$ $1.123^{\small+0.060}_{\small-0.056}$ $3.098^{\small+0.874}_{\small-0.905}$ $1.810^{\small+0.141}_{\small-0.209}$ $1.818^{\small+2.515}_{\small-1.375}$ $-0.366^{\small+0.681}_{\small-0.583}$ $11.353^{\small+0.279}_{\small-0.274}$ $-1.609^{\small+0.204}_{\small-0.189}$ $-1.582^{\small+0.320}_{\small-0.304}$ $0.618^{\small+0.154}_{\small-0.144}$ $3.273^{\small+0.648}_{\small-0.522}$
(3, 3) $0.162^{\small+0.044}_{\small-0.060}$ $1.410^{\small+0.098}_{\small-0.113}$ $2.826^{\small+5.573}_{\small-2.000}$ $1.210^{\small+0.177}_{\small-0.214}$ $1.753^{\small+6.226}_{\small-2.328}$ $0.145^{\small+0.315}_{\small-0.232}$ $11.375^{\small+0.278}_{\small-0.398}$ $-1.624^{\small+0.224}_{\small-0.227}$ $-1.814^{\small+0.301}_{\small-0.261}$ $0.395^{\small+0.066}_{\small-0.064}$ $3.463^{\small+0.602}_{\small-0.576}$
(4, 1) $0.102^{\small+0.097}_{\small-0.087}$ $1.437^{\small+0.087}_{\small-0.084}$ $19.613^{\small+2.758}_{\small-3.055}$ $0.853^{\small+0.306}_{\small-0.140}$ $2.980^{\small+4.147}_{\small-4.074}$ $0.407^{\small+0.706}_{\small-0.595}$ $10.656^{\small+0.310}_{\small-0.307}$ $-1.640^{\small+0.181}_{\small-0.139}$ $-2.141^{\small+0.296}_{\small-0.283}$ $0.522^{\small+0.153}_{\small-0.149}$ $4.467^{\small+0.589}_{\small-0.488}$
(4, 2) $0.460^{\small+0.102}_{\small-0.115}$ $1.481^{\small+0.105}_{\small-0.085}$ $9.834^{\small+3.727}_{\small-4.084}$ $1.580^{\small+0.292}_{\small-0.260}$ $3.187^{\small+3.487}_{\small-2.346}$ $-0.433^{\small+0.520}_{\small-0.393}$ $11.842^{\small+0.183}_{\small-0.174}$ $-1.826^{\small+0.077}_{\small-0.073}$ $-1.848^{\small+0.289}_{\small-0.283}$ $0.384^{\small+0.188}_{\small-0.181}$ $3.171^{\small+0.438}_{\small-0.383}$
(4, 3) $0.039^{\small+0.069}_{\small-0.069}$ $1.848^{\small+0.131}_{\small-0.163}$ $5.183^{\small+5.051}_{\small-4.289}$ $1.255^{\small+0.202}_{\small-0.153}$ $0.765^{\small+4.549}_{\small-3.520}$ $0.141^{\small+0.374}_{\small-0.266}$ $11.603^{\small+0.320}_{\small-0.428}$ $-1.727^{\small+0.209}_{\small-0.145}$ $-1.774^{\small+0.325}_{\small-0.282}$ $0.503^{\small+0.161}_{\small-0.136}$ $3.478^{\small+0.556}_{\small-0.417}$
@@ Acknowledgements
All analysis in this work was enabled greatly by the following software: NumPy (Van der Walt et al., 2011), SciPy (Virtanen et al., 2020), and Matplotlib (Hunter, 2007). We have also used the Astrophysics Data Service (ADS) and arXiv preprint repository extensively during this project and the writing of the paper.

### v3

@@ Appendix E Diagnostic tests
  • Random points: The mean tangential shear around random points tests the importance of geometrical and mask effects in the signal. Although our estimator of galaxy-galaxy lensing includes the subtraction of tangential shear measurement around random points, it is useful to check that this correction is small, especially for the bins with the highest signal. For this work we use a number of random points that is $30$ times the number of lenses in each bin. We note here, however, that we have tested different numbers of random points, both higher and lower, and we have verified that this number is sufficiently large for our purposes. See also Prat2022 for a discussion on random-point subtraction and its effects on the measured shear. The effect from including the random-point subtraction is shown on Figure 18, where we can see that it is a small effect, and within the error bars in all bins, as also indicated by the reduced $\chi^{2}$ values we obtain.
@@
  • Boost factors: Galaxies are clustered, and the number of sources around a lens can be larger than what we would expect from a uniform distribution if the lens and the source redshift distribution overlap. This is usually quantified by the boost factor (Sheldon_2004), $B(\theta)$, as estimator quantifying the excess in the number of sources around a lens with respect to randoms. This is a test only relevant for $\gamma_{t}$ and the results are plotted in Figure 18. As can be seen from the plot, the contribution from this effect can be large at small scales, especially when the bins are more overlapped in redshift. In our analysis we take the boost factors into account by correcting for it before carrying out the model fit (Equation 5.1). We also use the boost factors to guide our decision in the selection of bins and scales to use in the final analysis (Section 5.4).
@@ Appendix F Constraint on parameters
(1, 1) $0.427^{\small+0.144}_{\small-0.165}$ $1.378^{\small+0.079}_{\small-0.082}$ $14.077^{\small+2.551}_{\small-3.090}$ $0.202^{\small+0.073}_{\small-0.061}$ $9.411^{\small+3.051}_{\small-2.936}$ $0.419^{\small+0.228}_{\small-0.244}$ $11.624^{\small+0.101}_{\small-0.104}$ $1.068^{\small+0.172}_{\small-0.189}$ $2.007^{\small+0.329}_{\small-0.158}$ $0.208^{\small+0.192}_{\small-0.123}$ $5.883^{\small+0.473}_{\small-0.743}$
(1, 2) $0.329^{\small+0.069}_{\small-0.165}$ $0.628^{\small+0.046}_{\small-0.088}$ $7.483^{\small+6.117}_{\small-2.162}$ $1.739^{\small+0.205}_{\small-0.253}$ $10.133^{\small+3.321}_{\small-2.296}$ $0.141^{\small+0.380}_{\small-0.351}$ $11.168^{\small+0.157}_{\small-0.186}$ $1.736^{\small+0.095}_{\small-0.101}$ $1.940^{\small+0.329}_{\small-0.226}$ $0.555^{\small+0.198}_{\small-0.116}$ $3.807^{\small+0.576}_{\small-0.589}$
(1, 3) $0.142^{\small+0.127}_{\small-0.101}$ $0.791^{\small+0.046}_{\small-0.059}$ $7.552^{\small+1.024}_{\small-1.133}$ $1.999^{\small+0.280}_{\small-0.271}$ $7.295^{\small+3.541}_{\small-4.336}$ $0.767^{\small+0.278}_{\small-0.710}$ $11.287^{\small+0.094}_{\small-0.212}$ $1.700^{\small+0.093}_{\small-0.089}$ $2.016^{\small+0.262}_{\small-0.251}$ $0.679^{\small+0.131}_{\small-0.086}$ $3.874^{\small+0.658}_{\small-0.222}$
(1, 4) $0.146^{\small+0.015}_{\small-0.055}$ $1.087^{\small+0.004}_{\small-0.050}$ $8.193^{\small+0.027}_{\small-2.736}$ $1.554^{\small+0.290}_{\small-0.320}$ $5.592^{\small+1.800}_{\small-1.173}$ $0.284^{\small+0.485}_{\small-0.269}$ $11.147^{\small+0.437}_{\small-0.071}$ $1.893^{\small+0.082}_{\small-0.035}$ $1.977^{\small+0.589}_{\small-0.022}$ $0.855^{\small+0.089}_{\small-0.103}$ $5.414^{\small+0.130}_{\small-0.615}$
(2, 1) $0.492^{\small+0.060}_{\small-0.099}$ $1.232^{\small+0.027}_{\small-0.030}$ $18.918^{\small+1.351}_{\small-1.817}$ $1.289^{\small+0.129}_{\small-0.129}$ $0.963^{\small+3.079}_{\small-2.076}$ $1.446^{\small+0.443}_{\small-0.654}$ $10.519^{\small+0.114}_{\small-0.095}$ $1.968^{\small+0.088}_{\small-0.065}$ $1.220^{\small+0.308}_{\small-0.277}$ $0.796^{\small+0.062}_{\small-0.093}$ $4.652^{\small+0.260}_{\small-0.351}$
(2, 2) $0.414^{\small+0.092}_{\small-0.129}$ $1.230^{\small+0.053}_{\small-0.052}$ $9.170^{\small+2.976}_{\small-3.270}$ $1.147^{\small+0.200}_{\small-0.245}$ $0.864^{\small+3.844}_{\small-3.268}$ $1.524^{\small+0.620}_{\small-0.677}$ $10.609^{\small+0.156}_{\small-0.141}$ $1.744^{\small+0.125}_{\small-0.124}$ $1.571^{\small+0.288}_{\small-0.302}$ $0.623^{\small+0.130}_{\small-0.126}$ $3.948^{\small+0.522}_{\small-0.554}$
(2, 3) $0.442^{\small+0.109}_{\small-0.125}$ $1.232^{\small+0.051}_{\small-0.055}$ $8.058^{\small+2.303}_{\small-2.142}$ $1.541^{\small+0.182}_{\small-0.245}$ $1.034^{\small+2.307}_{\small-1.468}$ $0.348^{\small+0.678}_{\small-0.666}$ $11.113^{\small+0.180}_{\small-0.180}$ $1.864^{\small+0.157}_{\small-0.120}$ $1.577^{\small+0.304}_{\small-0.312}$ $0.635^{\small+0.102}_{\small-0.126}$ $4.322^{\small+0.527}_{\small-0.587}$
(2, 4) $0.487^{\small+0.250}_{\small-0.138}$ $1.173^{\small+0.037}_{\small-0.055}$ $4.053^{\small+1.528}_{\small-1.217}$ $1.560^{\small+0.360}_{\small-0.138}$ $0.300^{\small+1.785}_{\small-0.993}$ $0.710^{\small+0.371}_{\small-0.524}$ $11.857^{\small+0.385}_{\small-0.473}$ $1.870^{\small+0.228}_{\small-0.095}$ $2.056^{\small+0.426}_{\small-0.295}$ $0.599^{\small+0.170}_{\small-0.052}$ $3.740^{\small+0.932}_{\small-0.393}$
(3, 1) $0.236^{\small+0.111}_{\small-0.105}$ $1.165^{\small+0.063}_{\small-0.066}$ $7.050^{\small+2.638}_{\small-2.288}$ $1.004^{\small+0.197}_{\small-0.213}$ $2.622^{\small+3.662}_{\small-2.779}$ $1.527^{\small+0.683}_{\small-0.717}$ $10.746^{\small+0.164}_{\small-0.154}$ $1.545^{\small+0.140}_{\small-0.151}$ $1.730^{\small+0.312}_{\small-0.279}$ $0.484^{\small+0.152}_{\small-0.121}$ $3.441^{\small+0.578}_{\small-0.466}$
(3, 2) $0.259^{\small+0.128}_{\small-0.091}$ $1.123^{\small+0.060}_{\small-0.056}$ $3.098^{\small+0.874}_{\small-0.905}$ $10^{\small+0.141}_{\small-0.209} 1.818$$18^{\small+2.515}_{\small-1.375}$ $0.366^{\small+0.681}_{\small-0.583}$ $11.353^{\small+0.279}_{\small-0.274}$ $1.609^{\small+0.204}_{\small-0.189}$ $1.582^{\small+0.320}_{\small-0.304}$ $0.618^{\small+0.154}_{\small-0.144}$ $3.273^{\small+0.648}_{\small-0.522}$
(3, 3) $0.162^{\small+0.044}_{\small-0.060}$ $1.410^{\small+0.098}_{\small-0.113}$ $2.826^{\small+5.573}_{\small-2.000}$ $1.210^{\small+0.177}_{\small-0.214}$ $1.753^{\small+6.226}_{\small-2.328}$ $0.145^{\small+0.315}_{\small-0.232}$ $11.375^{\small+0.278}_{\small-0.398}$ $1.624^{\small+0.224}_{\small-0.227}$ $1.814^{\small+0.301}_{\small-0.261}$ $0.395^{\small+0.066}_{\small-0.064}$ $3.463^{\small+0.602}_{\small-0.576}$
(4, 1) $0.102^{\small+0.097}_{\small-0.087}$ $1.437^{\small+0.087}_{\small-0.084}$ $19.613^{\small+2.758}_{\small-3.055}$ $0.853^{\small+0.306}_{\small-0.140}$ $2.980^{\small+4.147}_{\small-4.074}$ $0.407^{\small+0.706}_{\small-0.595}$ $10.656^{\small+0.310}_{\small-0.307}$ $1.640^{\small+0.181}_{\small-0.139}$ $2.141^{\small+0.296}_{\small-0.283}$ $0.522^{\small+0.153}_{\small-0.149}$ $4.467^{\small+0.589}_{\small-0.488}$
(4, 2) $0.460^{\small+0.102}_{\small-0.115}$ $1.481^{\small+0.105}_{\small-0.085}$ $9.834^{\small+3.727}_{\small-4.084}$ $1.580^{\small+0.292}_{\small-0.260}$ $3.187^{\small+3.487}_{\small-2.346}$ $0.433^{\small+0.520}_{\small-0.393}$ $11.842^{\small+0.183}_{\small-0.174}$ $1.826^{\small+0.077}_{\small-0.073}$ $1.848^{\small+0.289}_{\small-0.283}$ $0.384^{\small+0.188}_{\small-0.181}$ $3.171^{\small+0.438}_{\small-0.383}$
(4, 3) $0.039^{\small+0.069}_{\small-0.069}$ $1.848^{\small+0.131}_{\small-0.163}$ $5.183^{\small+5.051}_{\small-4.289}$ $1.255^{\small+0.202}_{\small-0.153}$ $0.765^{\small+4.549}_{\small-3.520}$ $0.141^{\small+0.374}_{\small-0.266}$ $11.603^{\small+0.320}_{\small-0.428}$ $1.727^{\small+0.209}_{\small-0.145}$ $1.774^{\small+0.325}_{\small-0.282}$ $0.503^{\small+0.161}_{\small-0.136}$ $3.478^{\small+0.556}_{\small-0.417}$
@@ Acknowledgements
All analysis in this work was enabled greatly by the following software: NumPy (vanderWalt2011Numpy), SciPy (Virtanen2020Scipy), and Matplotlib (Hunter2007Matplotlib). We have also used the Astrophysics Data Service (ADS) and arXiv preprint repository extensively during this project and the writing of the paper.